# Probing thermonuclear bursts and X-ray reflection features in Aql X-1 during 2024 outburst

### v1

@@ Abstract
We report the broadband timing and spectral properties of the neutron star low-mass X-ray binary Aql X-1 during the 2024 outburst with NICER, NuSTAR, and Swift observatories. We detected six thermonuclear X-ray bursts during the NICER and NuSTAR observations, with the observed X-ray burst profiles exhibiting a strong energy dependence. The time-resolved burst spectra indicate the presence of soft excess during the burst, which can be modeled by using a variable persistent emission method ($f_{a}$ method), or the relxillNS reflection model. We found that the reflection model can contribute $\sim$20of total emission as observed during the NICER burst. The reflection and blackbody component fluxes are strongly correlated as observed during a burst. The excess emission is possible due to the enhanced mass accretion rate to the neutron star due to the Poynting-Rodertson drag and a fraction of burst emission may be reflected from the disk. The bursts did not show photospheric radius expansion during the peak. Moreover, we examined the burst-free accretion emission in the broadband range with NuSTAR, NICER, and Swift at two epochs of the outburst. The persistent emission showed X-ray reflection feature, which can be well modeled with the relativistic reflection model relxillCp. The inner disk radius (Rin) is found to be nearly 22 and 10 times $\rm R_{g}$ for two observations, respectively. Assuming that the inner disk is truncated at the magnetospheric radius, the magnetic field strength at the poles of the neutron star is estimated to be $(0.6-1.9)\times 10^{9}$ G.
@@ 1 Introduction
The unstable burning of accreted material such as hydrogen and/or helium on the surface of NS is responsible for a thermonuclear X-ray burst. During the burst, the lighter species transform into heavier elements through nuclear chain reactions (Lewin, van Paradijs & Taam, 1993; Strohmayer & Bildsten, 2003; Schatz & Rehm, 2006). The first thermonuclear X-ray burst, known as Type-I X-ray burst, was discovered in 1976 (Grindlay et al., 1976), which opened a new domain to study different properties of an NS. The Type-I X-ray bursts have been detected in over 100 sources to date [^1] https://personal.sron.nl/~jeanz/bursterlist.html (Liu et al., 2007; Galloway et al., 2020).
@@
Aquila X–1 (Aql X-1) is a NS LMXB discovered in 1965 (Friedman et al., 1967), showing a coherent X-ray pulsation (Casella et al., 2008) as well as burst oscillations at 550 Hz (Zhang et al., 1998; Galloway et al., 2008; Bilous & Watts, 2019). It shows frequent outbursts about once a year (Campana et al., 2014; Güngör, Güver, & Ekşi, 2014). Aql X-1 has a K-type donor at a distance of nearly 6 kpc (Mata Sánchez et al., 2017). Several efforts have been made to study thermonuclear bursts in Aql X-1 using data from various observatories (Güver et al., 2022; Keek et al., 2018; Galloway et al., 2008). Earlier, reflection from the accretion disk was also reported in the persistent spectra of Aql X-1 (King et al., 2016; Ludlam et al., 2017).
@@ 2.1 NuSTAR observation
The Nuclear Spectroscopic Telescope Array (NuSTAR) is an imaging X-ray telescope consisting of two identical detectors that are co-aligned and working in 3–79 keV energy band (, 2013). NuSTAR observed Aql X-1 twice under the Director’s Discretionary Time (DDT) proposal on MJD 60576.6 and MJD 60602.1 for effective exposures of around 20 ks each. The details of the observation are presented in Table LABEL:tab:log_table_burst. For data reduction, we used the standard NuSTAR data analysis software (NUSTARDAS v2.1.2) distributed with HEASOFT v 6.31.1 with the latest CALDB version 20240325. We used the nupipeline task to filter the event files. We used a circular region of radius 120 arcsec centered on the source location to extract the source events. A similar circular region was defined away from the centre of the source region to create background events. Finally, we used the nuproducts task to create the energy spectrum, the auxiliary response file, and the response matrix file for both detectors. The light curves are background corrected using the lcmath task. Further, we simultaneously modeled the energy spectra from both detectors to minimize the systematic effects.
@@ 3.2 Time-resolved burst spectroscopy
To probe the dynamic evolution of the X-ray bursts, time-resolved spectroscopy of each burst is performed to investigate the variation of different spectral parameters during the bursts. The pre-burst spectrum of each burst, either from NICER or NuSTAR observation, is modeled in XSPEC (Arnaud, 1996). The pre-burst NuSTAR spectra are obtained for a time bin of $\sim$300 s. The pre-burst spectra are modeled with a power law continuum along with a blackbody (bbodyrad) model. The TBABS (Wilms, Allen & McCray, 2000) model is used to account for the interstellar medium absorption. The best-fit pre-burst spectral fitting parameters are summarized in Table LABEL:tab:tab_preburst. The hydrogen column density ($N_{H}$) from the pre-burst spectra was $\sim 0.5\times 10^{22}$ cm-2.
To perform the time-resolved spectral study for each burst with NICER (0.5–10 keV), a total of 30 spectra (P1 to P30) are generated for each burst. The time bin of 1 s is used to create the NICER spectra up to the P15 segment, then the time bin is increased gradually to 2 s (for the segments P16–P25) and 4 s (for the segments P26–P30). To optimize the signal-to-noise ratio during the burst decay part, the time bin is extended from 2 s to 4 s. For the time-resolved spectral analysis with NuSTAR, each burst is divided into 8 segments (P1–P8). The burst rising part (P1) is for 5 s segment, the broad flat peak (P2) is for 8 s, and P3–P6 is for 8 s each. The bin time is increased to 15 s during the decay part.
@@ 3.2.1 Disk reflection during burst: burst-disk interaction
The deviation of the burst spectrum from a pure blackbody model can be probed with the disk reflection model. We used the latest relativistic reflection model relxillNS to investigate the reflection feature in the burst spectra of Aql X-1. In this model, the photoionized disk is illuminated by the blackbody spectrum from the neutron star (García et al., 2022). The NICER burst spectra can be modeled with tbabs $\times$(bbodyrad1+relxillNS+powerlaw+bbodyrad2), where bbodyrad1, relxillNS, and (powerlaw+bbodyrad2) represent the component of the burst, the reflection of burst photons from the accretion disk and persistent emission, respectively. During the fitting of burst spectra with the above reflection model, the preburst model parameters are fixed at the preburst values obtained during the earlier model fitting (Section 3.2). The relxillNS model parameters are: $q1$ and $q2$ are the inner and outer indices, $i$ is the disk inclination, $R_{in}$ and $R_{out}$ are the inner and outer disk radii, $a$ is the dimensionless spin parameter, $A_{Fe}$ is the iron abundance, $kT_{bb}$ represents the input blackbody temperature, log $\xi$ is the disk ionization parameter and $logN$ is the disk density. The emissivity profile of the accretion disk is defined as a broken power law with three governing parameters: index1 (q1), index2 (q2), and break radius ($\rm R_{br}$). Assuming a single emissivity profile ($r^{-q}$), The emissivity index is fixed at q1 = q2 = 3, which eliminates the significance of $R_{br}$. As the reflection model contains many parameters, it is challenging to constrain all parameters simultaneously in short intervals. Therefore, during the fitting, the temperature of the input blackbody spectrum is tied with the bbodyrad1 component, i=30, logN=17 cm-3, $A_{Fe}$=5, the inner and outer disk radii are fixed at $R_{ISCO}$, and 400 $R_{g}$, respectively. Initially, the disk ionization parameter (log $\xi$) was allowed to vary. We found that the variation in the ionization parameter is not significant during time-resolved spectroscopy. Therefore, we have fixed this parameter to a value of log $\xi$=3.2 erg cm s-1. The reflection fraction is fixed at -1, and the normalization of the relxillNS is allowed to vary during the burst. The F-test also indicates that the relxillNS model is statistically significant ($F$ value and probability of 20.5 and $2.5\times 10^{-10}$, respectively, during the peak). Therefore, we fitted the burst spectra with the reflection model for the NICER observation, for which the $f_{a}$ model was used. As the burst parameters showed a similar type of evolution during the NICER bursts, we showed the evolution of burst parameters with relxillNS model for the NICER burst-1 in Fig. 5. It can be seen that the contribution of the reflection component increases with the increase in the burst flux. During the burst peak, the flux (0.1–100 keV range) of the relxillNS and blackbody components reached a maximum value of $\sim 10^{-8}$ erg cm-2 s-1 and $\sim$ 3.5 ×$\times 10^{-8}$ erg cm-2 s-1, contributing about 20% and 74% towards the total flux, respectively. The blackbody temperature and flux are comparatively lower than the values obtained from the $f_{a}$ method.
@@ 3.3 Analysis of persistent spectra: Reflection
3.3.1 Modeling of the reflection features
We approach modeling the broadband energy spectra systematically. Before moving on to the complex models, we first applied simple single continuum models including the single temperature blackbody component (bbodyrad in XSPEC), the multi-temperature disk blackbody component (diskbb in XSPEC notations: Mitsuda et al. (1984); Makishima et al. (1986)), thermal Comptonization component (nthcomp in XSPEC notations: (Zdziarski et al., 1996; Zycki et al., 1999) and the power law model, modified by the TBabs model. Modeling with these single continuum models resulted in reduced chi-square ($\chi^{2}$) values $>$ 1.5, indicative of a poor data description. In the next step, we introduced moderately complex three-component models such as TBabs×(diskbb + nthComp) or TBabs×(diskbb + cutoffpl) to fit the broadband spectra. We observed a slight improvement in the reduced $\chi^{2}$. In the residuals (see Fig. 7), we detected X-ray reflection features including emission features in the 5–8 keV range, a photoelectric absorption dip around 10 keV, and a back-scattering Compton hump above 20 keV (Fabian et al., 1989, 2000; Miller et al., 2007). The reflection feature is evident irrespective of the choice of continuum components. Physically, these reflection features are most likely due to relativistic reflection from the accretion disk. The relxill/rexillCp is a self-consistent broadband reflection model from the relxill (v2.3: Garcia et al. (2014); Dauser et al. (2014)) model family, which describes relativistic reflection due to an incident cut-off power-law or nthcomp continuum. The broad-band spectrum shows a broad iron line around 6.4 keV followed by a Compton hump, which indicates the reflection feature. We further used a relativistic reflection model to probe this feature. The primary source is assumed with an emissivity given by Index1 (q1) and Index2 (q2). The reflection model is further defined by parameters such as the photon index of the illuminating radiation, the ionization parameter ($\xi$) at the disk’s surface, and the abundance of Fe relative to its solar value. We fitted the broadband spectra with the model TBabs×$\times$(nthcomp + diskbb + relxillCp) and found a significant improvement in the fit quality (reduced $\chi^{2}\sim$ 1). The best-fit broadband spectra are shown in Fig. 8 for two different NuSTAR observations. The bottom panel of Fig. 8 represents the residual, which indicates that the model combination can address all the features in the broadband spectra. The best-fit spectral parameters are summarized in Table 3.
[…]
@@ 3.3.1 Modeling of the reflection features
Figure 8: Best-fit broadband NICER and NuSTAR (obs-1) spectra (left panel) of Aql X-1 fitted with model M4 – tbabs×$\times$(nthcomp+diskbb+relxillCp). The right-hand side panel shows the broadband best-fit spectra for the Swift/XRT and NuSTAR (obs-2) spectra with the model M4. The bottom panel shows corresponding residuals.
@@ 4 Discussion
Using a large sample of bursts observed with NICER between 2017 July and 2021 April, the time-resolved burst spectral study of Aql X-1 reported a peak value of $f_{a}$ $<5$ for non-PRE bursts (Güver et al., 2022). However, in the case of the PRE bursts, the $f_{a}$ parameter increased to a maximum value of $\sim$10 (for details, see Fig. 4 (Güver et al., 2022)). There are several possible explanations for the deviation of the burst emission from a pure blackbody emission. The deviation is likely due to the atmospheric effect (Özel, 2013). The atmosphere model of NS predicts a broad X-ray spectrum rather than a pure blackbody (Suleimanov et al., 2011; Madej et al., 2004). The relativistic effect may also have an impact on broadening the features in the X-ray spectrum of pulsars (Bauböck et al., 2015). The $f_{a}$ method does not provide a detailed physical explanation of the increment of the pre-burst emission. However, the enhancement can be associated with the increase in the mass accretion rate to the NS due to the Poynting–Robertson drag (Walker et al., 1992; Worpel, Galloway & Price, 2013). The reflection of burst photons from the accretion disk is also anticipated (see, e.g., Güver et al. (2022)). Considering this, we attempted to investigate the contribution of disk reflection during the bursts. The soft excess emission during the burst can be modeled using a disk reflection model relxillNS. The reflection modeling approach provides a self-consistent and physically motivated explanation. This approach was adopted to model the time-resolved burst spectra for 4U 1820–30 (Keek et al., 2018; Jaisawal et al., 2024), 4U 1636–536 (Zhao et al., 2022), 4U 1730–22 (Lu et al., 2023), and SAX 1808.4–3658 (Bult et al., 2021). We found that the reflection component can contribute $\sim$20of the total burst emission during NICER burst. The blackbody temperature and blackbody flux obtained from the $f_{a}$ model are higher than those of the reflection model. During the burst, the flux of the reflection component is found to be strongly correlated with the blackbody flux as observed with NICER. Due to poor statistics in NuSTAR data, the reflection modeling does not result in any significant improvement over the $f_{a}$ modeling approach.
@@
where, $\dot{M}$ is the mass accretion rate as measured at the NS surface, $\zeta$ is the anisotropy factor, and $F$ is the pre-burst flux. The mass accretion rate is estimated using the average NuSTAR pre-burst flux (1–79 keV range) of 5.75$\sim$ 5.75$\times$ 5.75 ×10-9 erg cm2 s-1 and assuming the source distance of $\sim$6 kpc.
@@
The theoretical ignition models of hydrogen and helium-burning bursts can be utilized to probe the evolution of burst properties with accretion rates for different systems. For stable hot-CNO cycle hydrogen burning, the mass accretion rate is related to the Eddington mass accretion rate as $\dot{M}$geq 0.01 $\dot{M}_{Edd}$ (Cumming, 2004; Galloway & Cumming, 2006; Galloway et al., 2008). We estimated the mass accretion rate $\dot{M}$ $\simeq$ 0.18 $\dot{M}_{Edd}$ (assuming $\dot{M}_{Edd}=1.8\times 10^{-8}M_{\odot}yr^{-1}$), which is higher than the estimated rates by Cumming (2004) for stable CNO cycle hydrogen burning. Similarly, the preburst mass accretion rate is estimated from NICER data. The NICER preburst luminosity is found to vary between (1.2–2.8) × $\times\leavevmode\nobreak\ 10^{37}$ erg s(-1) for 3 observations. The corresponding mass accretion rate is estimated to be (1.5–3.6) × $\times\leavevmode\nobreak\ 10^{-9}$ $M_{\odot}yr^{-1}$ ($\sim$ (0.09–0.2) $\dot{M}_{Edd}$). At this level of mass accretion rate, the hydrogen burns stably, and the helium burning drives the thermal instability at a density of $\geq$ 105 gm cm-3. Based on the time to accumulate the critical column of fuel, the burst may be a “pure He” burst or “mixed H/He” burst Cumming (2004). However, the longer burst duration ($>$50 s) indicates that the burst may be powered by the H/He mixed fuel or H-rich fuel (Galloway & Keek, 2021).

### v2

@@ Abstract
We report the broadband timing and spectral properties of the neutron star low-mass X-ray binary Aql X-1 during the 2024 outburst with NICER, NuSTAR, and Swift observatories. We detected six thermonuclear X-ray bursts during the NICER and NuSTAR observations, with the observed X-ray burst profiles exhibiting a strong energy dependence. The time-resolved burst spectra indicate the presence of soft excess during the burst, which can be modeled by using a variable persistent emission method ($f_{a}$ method) or the relxillNS reflection model. We found that the reflection model can contribute $\sim$20of total emission as observed during the NICER burst. The reflection and blackbody component fluxes are strongly correlated, as observed during a burst. The excess emission is possible due to the enhanced mass accretion rate to the neutron star due to the Poynting-Rodertson drag, and a fraction of burst emission may be reflected from the disk. The bursts did not show photospheric radius expansion during the peak. Moreover, we examined the burst-free accretion emission in the broadband range with NuSTAR, NICER, and Swift at two epochs of the outburst. The persistent emission showed X-ray reflection features, which can be well modeled with the relativistic reflection model relxillCp. The inner disk radius (Rin) is found to be nearly 22 and 10 times $\rm R_{g}$ for two observations, respectively. Assuming that the inner disk is truncated at the magnetospheric radius, the magnetic field strength at the poles of the neutron star is estimated to be $(0.6-1.9)\times 10^{9}$ G.
@@ 1 Introduction
The unstable burning of accreted material, such as hydrogen and/or helium, on the surface of NS is responsible for a thermonuclear X-ray burst. During the burst, the lighter species transform into heavier elements through nuclear chain reactions (Lewin, van Paradijs & Taam, 1993; Strohmayer & Bildsten, 2003; Schatz & Rehm, 2006). The first thermonuclear X-ray burst, known as Type-I X-ray burst, was discovered in 1976 (Grindlay et al., 1976), which opened a new domain to study different properties of an NS. The Type-I X-ray bursts have been detected in over 100 sources to date [^1] https://personal.sron.nl/~jeanz/bursterlist.html (Liu et al., 2007; Galloway et al., 2020).
@@
Aquila X-1 (Aql X-1) is a NS LMXB discovered in 1965 (Friedman et al., 1967), showing a coherent X-ray pulsation (Casella et al., 2008) as well as burst oscillations at 550 Hz (Zhang et al., 1998; Galloway et al., 2008; Bilous & Watts, 2019). It shows frequent outbursts about once a year (Campana et al., 2014; Güngör, Güver, & Ekşi, 2014). Aql X-1 has a K-type donor at a distance of nearly 6 kpc (Mata Sánchez et al., 2017). Several efforts have been made to study thermonuclear bursts in Aql X-1 using data from various observatories (Güver et al., 2022; Keek et al., 2018; Galloway et al., 2008). Earlier, reflection from the accretion disk was also reported in the persistent spectra of Aql X-1 (King et al., 2016; Ludlam et al., 2017).
@@ 2.1 NuSTAR observation
The Nuclear Spectroscopic Telescope Array (NuSTAR) is an imaging X-ray telescope consisting of two identical detectors that are co-aligned and working in 3–79 keV energy band (, 2013). NuSTAR observed Aql X-1 twice under the Director’s Discretionary Time (DDT) proposal on MJD 60576.6 and MJD 60602.1 for effective exposures of around 20 ks each. The details of the observation are presented in Table LABEL:tab:log_table_burst. For data reduction, we used the standard NuSTAR data analysis software (NUSTARDAS v2.1.2) distributed with HEASOFT v 6.31.1 with the latest CALDB version 20240325. We used the nupipeline task to filter the event files. We used a circular region of radius 120 arcsec centered on the source location to extract the source events. A similar circular region was defined away from the centre of the source region to create background events. Finally, we used the nuproducts task to generate the energy spectrum, the auxiliary response file, and the response matrix file for both detectors. The light curves are background corrected using the lcmath task. Further, we simultaneously modeled the energy spectra from both detectors to minimize the systematic effects.
@@ 3.2 Time-resolved burst spectroscopy
To probe the dynamic evolution of the X-ray bursts, time-resolved spectroscopy of each burst is performed to investigate the variation of different spectral parameters during the bursts. The pre-burst spectrum of each burst, either from NICER or NuSTAR observation, is modeled in XSPEC (Arnaud, 1996). The pre-burst NuSTAR spectra are obtained for a time bin of $\sim$300 s. The pre-burst spectra are modeled with a power law continuum and a blackbody (bbodyrad) model. The TBABS (Wilms, Allen & McCray, 2000) model is used to account for the interstellar medium absorption. The best-fit pre-burst spectral fitting parameters are summarized in Table LABEL:tab:tab_preburst. The hydrogen column density ($N_{H}$) from the pre-burst spectra was $\sim 0.5\times 10^{22}$ cm-2.
To perform the time-resolved spectral study for each burst with NICER (0.5–10 keV), a total of 30 spectra (P1 to P30) are generated for each burst. The time bin of 1 s is used to create the NICER spectra up to the P15 segment, and then the time bin is increased gradually to 2 s (for the segments P16–P25) and 4 s (for the segments P26–P30). To optimize the signal-to-noise ratio during the burst decay part, the time bin is extended from 2 s to 4 s. For the time-resolved spectral analysis with NuSTAR, each burst is divided into 8 segments (P1–P8). The burst rising part (P1) is for 5 s segment, the broad flat peak (P2) is for 8 s, and P3–P6 is for 8 s each. The bin time is increased to 15 s during the decay part.
@@ 3.2.1 Disk reflection during burst: burst-disk interaction
The deviation of the burst spectrum from a pure blackbody model can be probed with the disk reflection model. We used the latest relativistic reflection model relxillNS to investigate the reflection feature in the burst spectra of Aql X-1. In this model, the photoionized disk is illuminated by the blackbody spectrum from the neutron star (García et al., 2022). The NICER burst spectra can be modeled with tbabs $\times$(bbodyrad1+relxillNS+powerlaw+bbodyrad2), where bbodyrad1, relxillNS, and (powerlaw+bbodyrad2) represent the component of the burst, the reflection of burst photons from the accretion disk and persistent emission, respectively. During the fitting of burst spectra with the above reflection model, the preburst model parameters are fixed at the preburst values obtained during the earlier model fitting (Section 3.2). The relxillNS model parameters are: $q1$ and $q2$ are the inner and outer indices, $i$ is the disk inclination, $R_{in}$ and $R_{out}$ are the inner and outer disk radii, $a$ is the dimensionless spin parameter, $A_{Fe}$ is the iron abundance, $kT_{bb}$ represents the input blackbody temperature, log $\xi$ is the disk ionization parameter and $logN$ is the disk density. The emissivity profile of the accretion disk is defined as a broken power law with three governing parameters: index1 (q1), index2 (q2), and break radius ($\rm R_{br}$). Assuming a single emissivity profile ($r^{-q}$), The emissivity index is fixed at q1 = q2 = 3, which eliminates the significance of $R_{br}$. As the reflection model contains many parameters, it is challenging to constrain all parameters simultaneously in short intervals. Therefore, during the fitting, the temperature of the input blackbody spectrum is tied with the bbodyrad1 component, i=30, logN=17 cm-3, $A_{Fe}$=5, the inner and outer disk radii are fixed at $\rm R_{ISCO}$, and 400 $R_{g}$, respectively. Initially, the disk ionization parameter (log $\xi$) was allowed to vary. We found that the variation in the ionization parameter is not significant during time-resolved spectroscopy. Therefore, we have fixed this parameter to a value of log $\xi$=3.2 erg cm s-1. The reflection fraction is fixed at -1, and the normalization of the relxillNS is allowed to vary during the burst. The F-test also indicates that the relxillNS model is statistically significant ($F$ value and probability of 20.5 and $2.5\times 10^{-10}$, respectively, during the peak). Therefore, we fitted the burst spectra with the reflection model for the NICER observation, for which the $f_{a}$ model was used. As the burst parameters showed a similar type of evolution during the NICER bursts, we showed the evolution of burst parameters with relxillNS model for the NICER burst-1 in Fig. 5. It can be seen that the contribution of the reflection component increases with the increase in the burst flux. During the burst peak, the flux (0.1–100 keV range) of the relxillNS and blackbody components reached a maximum value of $\sim 10^{-8}$ erg cm-2 s-1 and $\sim$ 3.5 ×$\times 10^{-8}$ erg cm-2 s-1, contributing about 20% and 74% towards the total flux, respectively. The blackbody temperature and flux are comparatively lower than the values obtained from the $f_{a}$ method.
@@ 3.3 Analysis of persistent spectra: Reflection
Figure 8: Best-fit broadband NICER and NuSTAR (obs-1) spectra (left panel) of Aql X-1 fitted with model M4 – tbabs×$\times$(nthcomp+diskbb+relxillCp). The right-hand side panel shows the broadband best-fit spectra for the Swift/XRT and NuSTAR (obs-2) spectra with the model M4. The bottom panel shows corresponding residuals. Data is rebinned to present in the plot.
3.3.1 Modeling of the reflection features
We approach modeling the broadband energy spectra systematically. Before moving on to the complex models, we first applied simple single continuum models including the single temperature blackbody component (bbodyrad in XSPEC), the multi-temperature disk blackbody component (diskbb in XSPEC notations: Mitsuda et al. (1984); Makishima et al. (1986)), thermal Comptonization component (nthcomp in XSPEC notations: (Zdziarski et al., 1996; Zycki et al., 1999) and the power law model, modified by the TBabs model. Modeling with these single continuum models resulted in reduced chi-square ($\chi^{2}$) values $>$ 1.5, indicative of a poor data description. In the next step, we introduced moderately complex three-component models such as TBabs×(diskbb + nthComp) or TBabs×(diskbb + cutoffpl) to fit the broadband spectra. We observed a slight improvement in the reduced $\chi^{2}$. In the residuals (see Fig. 7), we detected X-ray reflection features, including emission features in the 5–8 keV range, a photoelectric absorption dip around 10 keV, and a back-scattering Compton hump above 20 keV (Fabian et al., 1989, 2000; Miller et al., 2007). The reflection feature is evident irrespective of the choice of continuum components. Physically, these reflection features are most likely due to relativistic reflection from the accretion disk. The relxill/rexillCp is a self-consistent broadband reflection model from the relxill (v2.3: Garcia et al. (2014); Dauser et al. (2014)) model family, which describes relativistic reflection due to an incident cut-off power-law or nthcomp continuum. The broad-band spectrum shows a broad iron line around 6.4 keV followed by a Compton hump, which indicates the reflection feature. We further used a relativistic reflection model to probe this feature. The primary source is assumed to have an emissivity given by Index1 (q1) and Index2 (q2). The reflection model is further defined by parameters such as the photon index of the illuminating radiation, the ionization parameter ($\xi$) at the disk’s surface, and the abundance of Fe relative to its solar value. We fitted the broadband spectra with the model TBabs×$\times$(nthcomp + diskbb + relxillCp) and found a significant improvement in the fit quality (reduced $\chi^{2}\sim$ 1). The best-fit broadband spectra are shown in Fig. 8 for two different NuSTAR observations. The bottom panel of Fig. 8 represents the residual, which indicates that the model combination can address all the features in the broadband spectra. The best-fit spectral parameters are summarized in Table 3.
[…]
@@ 4 Discussion
Using a large sample of bursts observed with NICER between 2017 July and 2021 April, the time-resolved burst spectral study of Aql X-1 reported a peak value of $f_{a}$ $<5$ for non-PRE bursts (Güver et al., 2022). However, in the case of the PRE bursts, the $f_{a}$ parameter increased to a maximum value of $\sim$10 (for details, see Fig. 4 (Güver et al., 2022)). There are several possible explanations for the deviation of the burst emission from a pure blackbody emission. The deviation is likely due to the atmospheric effect (Özel, 2013). The atmosphere model of NS predicts a broad X-ray spectrum rather than a pure blackbody (Suleimanov et al., 2011; Madej et al., 2004). The relativistic effect may also have an impact on broadening the features in the X-ray spectrum of pulsars (Bauböck et al., 2015). The $f_{a}$ method does not provide a detailed physical explanation of the increment of the pre-burst emission. However, the enhancement can be associated with increased mass accretion rate to the NS due to the Poynting–Robertson drag (Walker et al., 1992; Worpel, Galloway & Price, 2013). The reflection of burst photons from the accretion disk is also anticipated (see, e.g., Güver et al. (2022)). Considering this, we attempted to investigate the contribution of disk reflection during the bursts. The soft excess emission during the burst can be modeled using a disk reflection model relxillNS. The reflection modeling approach provides a self-consistent and physically motivated explanation. This approach was adopted to model the time-resolved burst spectra for 4U 1820–30 (Keek et al., 2018; Jaisawal et al., 2024), 4U 1636–536 (Zhao et al., 2022), 4U 1730–22 (Lu et al., 2023), and SAX 1808.4–3658 (Bult et al., 2021). We found that the reflection component can contribute $\sim$20of the total burst emission during NICER burst. The blackbody temperature and blackbody flux obtained from the $f_{a}$ model are higher than those of the reflection model. During the burst, the flux of the reflection component is found to be strongly correlated with the blackbody flux as observed with NICER. Due to poor statistics in NuSTAR data, the reflection modeling does not result in any significant improvement over the $f_{a}$ modeling approach.
@@
where $\dot{M}$ is the mass accretion rate as measured at the NS surface, $\zeta$ is the anisotropy factor, and $F$ is the pre-burst flux. The mass accretion rate is estimated using the average NuSTAR pre-burst flux (1–79 keV range) of 5.75$\sim$ 5.75$\times$ 5.75 ×10-9 erg cm2 s-1 and assuming the source distance of $\sim$6 kpc.
@@
The theoretical ignition models of hydrogen and helium-burning bursts can be utilized to probe the evolution of burst properties with accretion rates for different systems. For stable hot-CNO cycle hydrogen burning, the mass accretion rate is related to the Eddington mass accretion rate as $\dot{M}$geq 0.01 $\dot{M}_{Edd}$ (Cumming, 2004; Galloway & Cumming, 2006; Galloway et al., 2008). We estimated the mass accretion rate $\dot{M}$ $\simeq$ 0.18 $\dot{M}_{Edd}$ (assuming $\dot{M}_{Edd}=1.8\times 10^{-8}M_{\odot}yr^{-1}$), which is higher than the estimated rates by Cumming (2004) for stable CNO cycle hydrogen burning. Similarly, the preburst mass accretion rate is estimated from NICER data. The NICER preburst luminosity is found to vary between (1.2–2.8) × $\times\leavevmode\nobreak\ 10^{37}$ erg s-1 for 3 observations. The corresponding mass accretion rate is estimated to be (1.5–3.6) × $\times\leavevmode\nobreak\ 10^{-9}$ $M_{\odot}yr^{-1}$ ($\sim$ (0.09–0.2) $\dot{M}_{Edd}$). At this level of mass accretion rate, the hydrogen burns stably, and the helium burning drives the thermal instability at a density of $\geq$ 105 gm cm-3. Based on the time to accumulate the critical column of fuel, the burst may be a “pure He” burst or “mixed H/He” burst Cumming (2004). However, the longer burst duration ($>$50 s) indicates that the burst may be powered by the H/He mixed fuel or H-rich fuel (Galloway & Keek, 2021).